# Rubber Friction: Theory, Mechanisms, and Challenges

### v2

@@ 1 Introduction
Sec. 9 reviews the similarities and differences between friction on ice and snow. While the mechanisms of friction on ice have been widely studied due to their relevance to winter mobility, rubber–snow interactions remain less well understood. The temperature dependence of the friction behavior on ice as well as the water content dependence on snow are discussed. Experimental findings are compared with theoretical predictions, highlighting the role of surface premelting, capillary forces, and viscoelastic dissipation.
@@ 2 Viscoelastic modulus
When deformed slowly at temperatures well below $T_{\rm g}$, rubber is in the glassy state, where molecular motion is limited. At temperatures well above $T_{\rm g}$, it is in the rubbery state with high chain mobility, and would be a (high-viscosity) liquid if not for the crosslinks.
@@
with typical values $C\approx 17.44$ and $T_{1}\approx 51.6\ {\rm K}$. However, for most rubber compounds, this relation is valid only within a limited temperature range. This is illustrated in Fig. 3, which shows the shift factor $a_{T}$ for a carbon-filled styrene-butadiene rubber (SBR) compound. It should also be noted that the frequency–temperature shifting procedure used to construct the master curve is strictly valid only at very small strains, where stress and strain are linearly related. In the linear response region, ${\rm Re}E$ and ${\rm Im}E$ are not independent functions: ${\rm Re}E$ can be obtained from ${\rm Im}E$ using a Kramers–Kronig relation. Hence, if shifting produces a smooth mastercurve for, e.g., ${\rm Im}E$, the same shift function $a_{T}$ will result in a smooth mastercurve for ${\rm Re}E$. This is not true at larger strains, where there is no simple relation between the real and imaginary parts of the effective modulus.
@@
In Fig. 4, we show $\eta_{\rm R}$ and $\eta_{\rm I}$ for SBR compound with carbon-filler. The function $\eta_{\rm R}(\epsilon)$ decreases monotonically with increasing strain. In contrast, $\eta_{\rm I}(\epsilon)$ first increases and reaches a maximum around $\epsilon\approx 0.01$, then decreases continuously over the strain range studied. For very large strain, where the polymer chains are fully stretched, $\eta(\epsilon)$ increases rapidly, but this large strain region is not important in most applications.
@@ 5 Rubber friction on rough surfaces
$\mu=\mu_{\rm visc}+\mu_{\rm con}$ (12)
Experiments have shown that for sliding on randomly rough surfaces, both the viscoelastic contribution $\mu_{\rm visc}$ and the contact-area contribution $\mu_{\rm con}$ are independent of $F_{\rm N}$, unless the normal force is large enough to approach full contact, or the sliding speed is high enough for frictional heating to become significant, or the surfaces are so smooth that adhesion manifest itself on the macroscopic scale as a pull-off force.
@@
Figure 12: The frictional shear stress in the area of real contact [see Fig. 10(b)] can arise from (a)-(c) binding–stretching–detachment cycles of rubber molecules at the sliding interface, or (d) from shearing of thin contamination films located between the substrate and the rubber.
Figure 13: Calculated frictional shear stress (a) and the fraction $N/N_{0}$ of the rubber surface in the sliding state (b), shown as a function of the logarithm of the sliding speed. $D$ is the characteristic size of the interfacial units that undergo stick–slip–stick motion. Results are for styrene–butadiene (SB) rubber. Adapted from Ref. [21].
@@
We first consider dry and clean surfaces. In sliding contact, rubber molecules and substrate atoms interact as illustrated in Fig. 12(a)–(c). In many cases, the interfacial interactions are weak, such as van der Waals forces. During stationary contact, the rubber chains at the interface adjust to the substrate potential in order to minimize the free energy. This bond formation may involve overcoming energy barriers and therefore does not occur instantaneously, but requires a finite relaxation time.
During sliding at low velocities, thermal fluctuations assist in formation and breaking of the rubber–substrate bonds, resulting in a friction force that approaches zero as the sliding velocity tends to zero. At high velocities, there is not enough time for the rubber molecules to adjust to the substrate potential. As a result, for high sliding speeds the rubber interface effectively “floats” above the substrate in an incommensurate-like state relative to the corrugated potential. In this regime, the frictional shear stress is also small.
Therefore, the frictional shear stress as a function of sliding speed must exhibit a maximum at some intermediate velocity $v^{*}$. This friction mechanism was first studied in a simplified model by Schallamach [16], and later extended by Chernyak and Leonov [20], and in a more realistic model by Persson and Volokitin [21].
The actual elementary stick-slip units are not likely to be individual molecular segments, but rather small patches of molecular segments with a characteristic size $D$. In Fig. 13, we show the results of calculations for $D=30$ and $40\ {\rm nm}$. Fig. 13(a) shows the frictional shear stress, and Fig. 13(b) shows the fraction $N/N_{0}$ of the rubber surface in the sliding state, both as functions of the logarithm of the sliding speed [21].
@@
Opening and closing crack propagation governs the friction behavior when soft rubber slides on a substrate with very smooth surfaces on both side. This is illustrated in Fig. 15, where a smooth rubber block slides on a flat substrate with negligible roughness. In this case, the rubber adheres to the substrate, and no direct slip occurs at the interface. Instead, localized detachment zones form at the leading edge and propagate opposite to the sliding direction. These are known as Schallamach waves.
Schallamach wave motion can be understood by analogy with the motion of a ruck on a carpet: one can either pull the entire carpet or create a local deformation and move it forward. Due to adhesion hysteresis, more energy is required to break the bonds at the opening crack than is recovered at the closing crack. This results in a friction force that can be estimated as follows.
@@
There are $N=L_{x}/\lambda$ waves, each of length $L_{y}$, and the energy dissipated is $NL_{y}\gamma_{\rm eff}v\Delta t$, where $\gamma_{\rm eff}=\gamma_{\rm open}-\gamma_{\rm close}\approx\gamma_{\rm open}$ is the effective adhesion energy. Equating the two gives
@@ 5.1 Analytical theory of rubber friction on rough surfaces
$\mu=\mu_{\rm visc}+\frac{\tau_{\rm f}A}{\sigma_{0}A_{0}}$ (16)
where $A$ is the area of real contact and $\tau_{\rm f}$ is the frictional shear stress acting within the contact area.
The Persson contact mechanics theory predicts the viscoelastic contribution $\mu_{\rm visc}$ and the area of real contact $A$. However, the dependence of the frictional shear stress $\tau_{\rm f}$ on sliding velocity and temperature must be determined using other theoretical approaches or extracted from experimental data.
@@
We now consider the contribution to the friction force arising from the area of real contact $A$. In sliding contact, rubber molecules interact with substrate atoms as illustrated in Fig. 10(a)–(c).
For this case, theory [21] predicts that the frictional shear stress exhibits a Gaussian-like dependence on the logarithm of the sliding speed. The curve has a full width at half maximum of $\sim 4$ (or more) frequency decades, and, at room temperature, is centered around a characteristic sliding speed typically on the order of $v^{*}\approx 1\ {\rm cm/s}$.
In Ref. [27], we found that using the following empirical shear stress law $\tau_{\rm f}(v,T)$, which is similar to the theoretical prediction [21], resulted in good agreement between theoretical results and experimental measurements:
$\tau_{\rm f}(v)\approx\tau_{\rm f0}\,\exp\left(-c\left[\log_{10}\left(\frac{v}{v^{*}}\right)\right]^{2}\right)$ (21)
where for passenger car tread rubber, typically $c\approx 0.17$, $\tau_{\rm f0}\approx 4$$8\ {\rm MPa}$, and the reference sliding speed is $v^{*}\approx 1\ {\rm cm/s}$ at $T=20^{\circ}{\rm C}$.
The full width at half maximum of $\tau_{\rm f}(v)$ as a function of $\log_{10}v$ is given by $\Gamma=2({\rm ln}2/c)^{1/2}\approx 4$.
@@
where $\epsilon\approx 1\ {\rm eV}$ is an activation energy, and $k_{\rm B}$ is the Boltzmann constant.
@@
[…]
As with the friction force, both the frictional shear stress acting in the area of real contact (adhesive contribution) and the viscoelastic deformations contribute to frictional heating. We denote the corresponding friction coefficients as $\mu_{\rm con}$ and $\mu_{\rm visc}$, respectively. The adhesive contribution results in energy dissipation within a nanometer-thick surface layer.
@@
In the following analysis, we assume that the macroasperities have a diameter of $d\approx 0.1\ {\rm mm}$, which is typical for rubber in contact with rough road surfaces.
Consider first the adhesive contribution. The frictional energy dissipated per unit area and unit time in the asperity contact regions is given by $q=\mu_{\rm con}\sigma v$, where $\sigma=F_{\rm N}/A$ is the normal contact stress. The asperity contact time is $\tau=d/v$, and the energy dissipated per unit volume is $Q=q\tau/h=qd/hv=\mu_{\rm con}\sigma d/h$, where $h\approx 1\ {\rm nm}$ is the thickness of the region at the surface where the adhesive frictional energy is deposited.
@@
where $D=\kappa_{\rm th}/\rho C_{V}$ is the thermal diffusivity, and $\rho$, $C_{V}$, and $\kappa_{\rm th}$ are the rubber mass density, heat capacity, and thermal conductivity, respectively.
@@
Note that $J_{d}=J_{h}^{2}$ when $v=Dd/h^{2}$, which, for the adhesive contribution where $h\approx 1\ {\rm nm}$, corresponds to an extremely high velocity in most cases. For example, if $d=0.1\ {\rm mm}$, then $v\approx 10^{7}\ {\rm m/s}$. Hence, for the adhesive contribution, the $J_{h}^{2}$ term can always be neglected.
In contrast, for the viscoelastic contribution, energy is dissipated in a volume element extending a distance $h\approx d$ into the rubber, so that $v=Dd/h^{2}\approx D/d\approx 1\ {\rm mm/s}$ for $d=0.1\ {\rm mm}$. Thus, for the viscoelastic case, the $J_{h}^{2}$ term cannot be neglected. However, in this case a more detailed analysis (as presented in Ref. [22]) is required, since viscoelastic energy dissipation decays rapidly with depth into the rubber.
Since $J_{d}=1$ corresponds to the sliding speed $v=D/d\approx 10^{-3}\ {\rm m/s}$ (using $d=0.1\ {\rm mm}$), and since for this velocity $\Delta T\approx qd/\kappa_{\rm th}=\mu\sigma vd/\kappa_{\rm th}\approx 0.1$$1\ {\rm K}$ for $\sigma\approx 1$$10\ {\rm MPa}$ (see below), we conclude that frictional heating is negligible at sliding speeds below $v\approx 10^{-3}\ {\rm m/s}$.
@@
$\Delta T\approx{qd\over\kappa_{\rm th}}(0.39J_{d})^{-1/2}\approx 1.6{\mu_{\rm con}\sigma\over\kappa_{\rm th}}(vDd)^{1/2}$ (27)
@@
Figure 17: Flash temperature as a function of sliding speed, estimated using (27) with typical parameter values: $D=10^{-7}\ {\rm m^{2}/s}$, $\kappa_{\rm th}=0.5\ {\rm W/(K\cdot m)}$, $d=0.1\ {\rm mm}$, $\sigma=10^{7}\ {\rm Pa}$, and $\mu_{\rm con}=0.5$.
@@
In our analysis of rubber friction on road surfaces, we have found that including surface roughness down to length scales where the cumulative rms slope $\xi$ is of order 1 yields results in good agreement with experimental data. This typically corresponds to a cut-off wavenumber of $q_{1}=2\pi/\lambda_{\rm min}\approx 10^{6}\ {\rm m}^{-1}$.
@@
Even if the exact value of $\lambda_{\rm min}$ is not known, the total friction coefficient can still be expressed as in (12), where the contact area contribution is written as $\mu_{\rm con}=\tau_{\rm eff}A/\sigma_{0}A_{0}$. Here, $A$ is the real contact area, calculated by including all surface roughness down to the assumed cut-off length. In this case the shear stress is an effective shear stress which could have contributions both from viscoelasticity and from the adhesive interaction in the area of real contact.
@@
Thus, in some studies, $A$ has been chosen as the nominal contact area $A_{0}$, in which case $\mu_{\rm visc}$ arises solely from macroscopic deformations of the contact region. This approach was employed in Ref. [56], where a rigid sphere was slid on human skin, a viscoelastic material. In that case, the nominal contact was Hertzian-like, and the macroscopic contact area had a distorted circular shape. Under such conditions, all asperity-induced viscoelastic energy dissipation is effectively captured within the definition of $\tau_{\rm eff}$.
@@
Another important issue concerns the role of adhesion. Here, we are not referring to the adhesive interaction between rubber molecules and the substrate within the real contact area, which gives rise to the frictional shear stress $\tau_{\rm f}$. Instead, we refer to two distinct adhesion-related mechanisms: (i) the increase in real contact area due to adhesion, and (ii) the energy dissipation from opening cracks at the edges of asperity contacts, occurring at length scales where adhesion is relevant.
@@
The contribution of adhesion to the enhancement of contact area was investigated by Plagge and Hentschke [52], while the contribution from opening cracks was studied by Le Gal et al. [57] and Carbone et al. [38]. These mechanisms may significantly affect friction in very soft rubber compounds sliding on smooth surfaces. However, for rough surfaces with roughness on many length scales, such as road surfaces, experimental evidence suggests that this form of adhesion is often negligible.

### v3

@@ 1 Introduction
Sec. 9 reviews the similarities and differences between friction on ice and snow. While the mechanisms of friction on ice have been widely studied due to their relevance to winter mobility, rubber-snow interactions remain less well understood. The temperature dependence of the friction behavior on ice as well as the water content dependence on snow are discussed. Experimental findings are compared with theoretical predictions, highlighting the role of surface premelting, capillary forces, and viscoelastic dissipation.
@@ 2 Viscoelastic modulus
When deformed at temperatures well below $T_{\rm g}$, rubber is in the glassy state, where molecular motion is limited. At temperatures well above $T_{\rm g}$, it is in the rubbery state with high chain mobility, and would be a (high-viscosity) liquid if not for the crosslinks.
@@
with typical values $C\approx 17.44$ and $T_{1}\approx 51.6\ {\rm K}$. However, for most rubber compounds, this relation is valid only within a limited temperature range. This is illustrated in Fig. 3, which shows the shift factor $a_{T}$ for a carbon-filled styrene-butadiene rubber (SBR) compound. It should also be noted that the frequency-temperature shifting procedure used to construct the master curve is strictly valid only at very small strains, where stress and strain are linearly related. In the linear response region, ${\rm Re}E$ and ${\rm Im}E$ are not independent functions: ${\rm Re}E$ can be obtained from ${\rm Im}E$ using a Kramers-Kronig relation. Hence, if shifting produces a smooth mastercurve for, e.g., ${\rm Im}E$, the same shift function $a_{T}$ will result in a smooth mastercurve for ${\rm Re}E$. This is not true at larger strains, where there is no simple relation between the real and imaginary parts of the effective modulus.
@@
In Fig. 4, we show $\eta_{\rm R}$ and $\eta_{\rm I}$ for a SBR compound with carbon-filler. The function $\eta_{\rm R}(\epsilon)$ decreases monotonically with increasing strain. In contrast, $\eta_{\rm I}(\epsilon)$ first increases and reaches a maximum around $\epsilon\approx 0.01$, then decreases continuously over the strain range studied. For very large strain, where the polymer chains are fully stretched, $\eta_{\rm R}(\epsilon)$ increases rapidly, but this large strain region is not important in most applications.
@@ 5 Rubber friction on rough surfaces
$\mu=\mu_{\rm visc}+\mu_{\rm cont}$ (12)
Experiments have shown that for sliding on randomly rough surfaces, both the viscoelastic contribution $\mu_{\rm visc}$ and the contact-area contribution $\mu_{\rm cont}$ are independent of $F_{\rm N}$, unless the normal force is large enough to approach full contact, or the sliding speed is high enough for frictional heating to become significant, or the surfaces are so smooth that adhesion manifest itself on the macroscopic scale as a pull-off force.
@@
Figure 12: The frictional shear stress in the area of real contact [see Fig. 10(b)] can arise from (a)-(c) binding-stretching-detachment cycles of rubber molecules at the sliding interface, or (d) from shearing of thin contamination films located between the substrate and the rubber.
Figure 13: Calculated frictional shear stress (a) and the fraction $N/N_{0}$ of the rubber surface in the sliding state (b), shown as a function of the logarithm of the sliding speed. $D$ is the characteristic size of the interfacial units that undergo stick-slip-stick motion. Results are for styrene-butadiene (SB) rubber. Adapted from Ref. [21].
@@
We first consider dry and clean surfaces. In sliding contact, rubber molecules and substrate atoms interact as illustrated in Fig. 12(a)-(c). In many cases, the interfacial interactions are weak, such as van der Waals forces. During stationary contact, the rubber chains at the interface adjust to the substrate potential in order to minimize the free energy. This bond formation may involve overcoming energy barriers and therefore does not occur instantaneously, but requires a finite relaxation time.
During sliding at low velocities, thermal fluctuations assist in formation and breaking of the rubber-substrate bonds, resulting in a friction force that approaches zero as the sliding velocity tends to zero. At high velocities, there is not enough time for the rubber molecules to adjust to the substrate potential. As a result, for high sliding speeds the rubber interface effectively “floats” above the substrate in an incommensurate-like state relative to the corrugated potential. In this regime, the frictional shear stress is also small.
Therefore, the frictional shear stress as a function of sliding speed must exhibit a maximum at some intermediate velocity $v^{*}$. This friction mechanism was first studied in a simplified model by Schallamach [16], and later extended by Chernyak and Leonov [20], and in a more realistic model by Persson and Volokitin [21]. Because of its conceptual importance, we present in Appendix A a derivation of the Schallamach result using an approach that differs slightly from the one originally used by Schallamach.
The actual elementary stick-slip units are not likely to be individual molecular segments, but rather small patches of molecular segments, which we refer too as stress domains, with a characteristic size $D$. In Fig. 13, we show the results of calculations (using the model of Persson and Volokitin) for $D=30$ and $40\ {\rm nm}$. Fig. 13(a) shows the frictional shear stress, and Fig. 13(b) shows the fraction $N/N_{0}$ of the rubber surface in the sliding state, both as functions of the logarithm of the sliding speed [21].
In the simple Schallamach model (see Appendix A), the frictional shear stress is given by $\sigma_{\rm f}=\sigma^{*}f(v/v^{*})$, where $\sigma^{*}$ is the maximum shear stress (with $D=2R$):
$\sigma^{*}\approx\left({2Ek_{\rm B}T\over R^{3}}\right)^{1/2}$ (14)
and the reference velocity:
$v^{*}=\left({2k_{\rm B}T\over ER\tau_{0}^{2}}\right)^{1/2}$ (15)
Here, $k_{\rm B}$ is the Boltzmann constant, $T$ the absolute temperature, $E$ the elastic modulus, and $\tau_{0}$ the average time a stress domain spends in the detached state.
The function $f(v/v^{*})$ depends on $\Delta E=\epsilon_{1}-\epsilon_{0}$, where $\epsilon_{1}$ is the energetic barrier to break the rubber-substrate bond in a stress domain, and $\epsilon_{0}$ the energetic barrier for forming the bond. For $\Delta E/k_{\rm B}T\approx 3$, $f(v/v^{*})$ is approximately a Gaussian function of ${\rm log}_{10}v$, with the full width at half maximum of about 2 velocity decades.
The values for $\sigma^{*}$ and $v^{*}$ are consistent with experimental observations (see Appendix A). Experiments show that the full width at half maximum of $\sigma_{\rm f}(v)$ is about 4 velocity decades, which is larger than predicted by the Schallamach model, but similar to the prediction by the Persson-Volokitin theory (see Fig. 13).
@@
In classical theories of interfacial crack propagation, it is typically assumed that the stress field extends far from the crack tip into the contact region and where it remains in a relaxed state. However, this assumption does not hold in situations where an indenter, such as a rigid sphere, is rapidly brought into and lifted out of contact with a viscoelastic solid. A similar condition applies to most sliding friction scenarios, where the rubber at a given surface point is initially nearly undeformed but undergoes rapid deformation as it enters and exits contact with substrate asperities, particularly at high sliding speeds. In such cases, the stress remains confined to the immediate vicinity of the contact region and does not penetrate deeply into the bulk. The influence of time-dependent deformation on interfacial crack propagation has been discussed qualitatively in Ref. [43] and quantitatively in Ref. [93] and [44].
Opening and closing crack propagation govern the friction behavior when soft rubber slides on a substrate, where both the rubber and the substrate surface are very smooth. This is illustrated in Fig. 15, where a smooth rubber block slides on a flat substrate with negligible roughness. In this case, the rubber adheres to the substrate, and no direct slip occurs at the interface. Instead, localized detachment zones form at the leading edge and propagate opposite to the sliding direction. These are known as Schallamach waves.
Schallamach wave motion can be understood by analogy with the motion of a rug on a carpet: one can either pull the entire carpet or create a local deformation and move it forward. Due to adhesion hysteresis, more energy is required to break the bonds at the opening crack than is recovered at the closing crack. This results in a friction force that can be estimated as follows.
@@
There are $N=L_{x}/\lambda$ waves, each of length $L_{y}$, and the energy dissipated is $NL_{y}\gamma_{\rm eff}v\Delta t$, where $\gamma_{\rm eff}=\gamma_{\rm open}-\gamma_{\rm close}\approx\gamma_{\rm open}$ is the effective adhesion energy. Equating this expression for the energy dissipation with $F_{\rm f}v_{0}\Delta t$ gives
@@ 5.1 Analytical theory of rubber friction on rough surfaces
$\mu=\mu_{\rm visc}+\frac{\sigma_{\rm f}A}{\sigma_{0}A_{0}}$
where $A$ is the area of real contact and $\sigma_{\rm f}$ is the frictional shear stress acting within the contact area.
The Persson contact mechanics theory predicts the viscoelastic contribution $\mu_{\rm visc}$ and the area of real contact $A$. However, the dependence of the frictional shear stress $\sigma_{\rm f}$ on sliding velocity and temperature must be determined using other theoretical approaches or extracted from experimental data.
@@
The general theory also includes temperature effects, both the flash temperature and the gradual heating of the sliding rubber block with increasing sliding speed [22]. The theory can also be applied to layered materials, e.g., for a rubber film on another solid. For a layered material consisting of two materials $\bf 0$} and $\bf 1$}, with 0 forming a film of thickness $d$ on top of a semi-infinite 1, one needs to replace $E^{*}(\omega)$ in (17)-(19) with $E^{*}_{0}(\omega)/K(\omega)$, where
$K=\frac{1+4mqde^{-2qd}-mne^{-4qd}}{1-(m+n+4mq^{2}d^{2})e^{-2qd}+mne^{-4qd}}$
and
$m=\frac{\mu_{0}/\mu_{1}-1}{\mu_{0}/\mu_{1}+3-4\nu_{0}},\ \ \ \ \ \ n=1-\frac{4(1-\nu_{0})}{1+(\mu_{0}/\mu_{1})(3-4\nu_{1})}$
The shear modulus $\mu_{0}=E_{0}/2(1+\nu_{0})$, and similarly for $\mu_{1}$. Note that $K$ is dimensionless and depends only on $qd$, $\nu_{0}$, $\nu_{1}$, and $E_{0}/E_{1}$. Note also that $K=1$ for $d=\infty$, and $K=E_{0}^{*}/E_{1}^{*}$ for $d=0$.
As an example, consider a thin rubber film (thickness $d$) on a rigid substrate sliding on a concrete surface. Fig. 16 shows the viscoelastic contribution $\mu_{\rm visc}$ to the sliding friction coefficient, and the relative contact area $A/A_{0}$ as a function of the sliding speed for the film thicknesses $d=0.1$, $0.3$, and $1\ {\rm mm}$, and for $d=\infty$. The concrete surface is relatively smooth, with an rms roughness $h_{\rm rms}\approx 0.05\ {\rm mm}$, which is considerably smaller than all the film thicknesses used. However, $h_{\rm rms}$ does not determine the depth of the deformation field into the rubber; this is determined by the width of the asperity contact regions. The surface used has a small-wavenumber cut-off $q_{0}\approx 615\ {\rm m^{-1}}$, corresponding to a wavelength $\lambda_{0}=2\pi/q_{0}\approx 10\ {\rm mm}$, and the width of the largest asperity contact regions will be some fraction of this. This explains why the friction drops already for $d<1\ {\rm mm}$.
Figure 16: The viscoelastic contribution $\mu_{\rm visc}$ to the sliding friction coefficient, and the relative contact area $A/A_{0}$, as functions of the sliding speed for a flat rigid solid covered with a tread rubber film of thickness $d$ sliding on a concrete surface. The concrete surface is relatively smooth, with an rms roughness of $0.05\ {\rm mm}$, and the calculation includes all roughness components down to a cut-off wavenumber of $\approx 10^{6}\ {\rm m^{-1}}$, where the rms slope is $\approx 1$. Frictional heating is not included in the calculation. The nominal contact pressure is $\sigma_{0}=0.2\ {\rm MPa}$.
@@
We now consider the contribution to the friction force arising from the area of real contact $A$. In sliding contact, rubber molecules interact with substrate atoms as illustrated in Fig. 10(a)-(c).
For this case, theory [21] (see also Appendix A) predicts that the frictional shear stress exhibits a Gaussian-like dependence on the logarithm of the sliding speed. The curve has a full width at half maximum of $\sim 4$ (or more) frequency decades, and, at room temperature, is centered around a characteristic sliding speed typically on the order of $v^{*}\approx 1\ {\rm cm/s}$.
In Ref. [27], we found that using the following empirical shear stress law $\sigma_{\rm f}(v,T)$, which is similar to the theoretical prediction [21], resulted in good agreement between theoretical results and experimental measurements:
$\sigma_{\rm f}(v)\approx\sigma_{\rm f0}\,\exp\left(-c\left[\log_{10}\left(\frac{v}{v^{*}}\right)\right]^{2}\right)$ (21)
where for passenger car tread rubber, typically $c\approx 0.17$, $\sigma_{\rm f0}\approx 4$-$8\ {\rm MPa}$, and the reference sliding speed is $v^{*}\approx 1\ {\rm cm/s}$ at $T=20^{\circ}{\rm C}$.
The full width at half maximum of $\sigma_{\rm f}(v)$ as a function of $\log_{10}v$ is given by $\Gamma=2({\rm ln}2/c)^{1/2}\approx 4$.
@@
where $\epsilon\approx 1\ {\rm eV}$ is an activation energy, and $k_{\rm B}$ is the Boltzmann constant. This temperature dependence is slightly weaker than that predicted by the bulk viscoelastic shift factor and may reflect a higher mobility of the polymer chains at the rubber surface compared with the bulk.
@@
At low sliding speeds $v$ (typically below 11-$10\ {\rm mm/s}$), thermal diffusion significantly reduces the flash temperature. However, at higher sliding speeds, the temperature increase $\Delta T$ is large, and will shifts both the viscoelastic and adhesive friction contributions toward higher sliding speeds.
As with the friction force, both the frictional shear stress acting in the area of real contact (adhesive contribution) and the viscoelastic deformations contribute to frictional heating. We denote the corresponding friction coefficients as $\mu_{\rm cont}$ and $\mu_{\rm visc}$, respectively. The adhesive contribution results in energy dissipation within a nanometer-thick surface layer.
@@
In the following analysis, we assume that the macroasperities have a diameter of $d\approx 1\ {\rm mm}$, which is typical for rubber in contact with rough road surfaces.
Consider first the adhesive contribution. The frictional energy dissipated per unit area and unit time in the asperity contact regions is given by $q=\mu_{\rm cont}\sigma v$, where $\sigma=F_{\rm N}/A$ is the normal contact stress. The asperity contact time is $\tau=d/v$, and the energy dissipated per unit volume is $Q=q\tau/h=qd/hv=\mu_{\rm cont}\sigma d/h$, where $h\approx 1\ {\rm nm}$ is the thickness of the region at the surface where the adhesive frictional energy is deposited.
@@
where $D=\kappa_{\rm th}/\rho C_{V}$ is the thermal diffusivity, and $\rho$, $C_{V}$, and $\kappa_{\rm th}$ are the rubber mass density, heat capacity, and thermal conductivity, respectively. For rubber with carbon fillers, typical values are $\rho\approx 1500\ {\rm kg/m^{3}}$, $C_{V}\approx 1500\ {\rm J/(kg\,K)}$, and $\kappa\approx 0.3\ {\rm W/(m\,K)}$, giving a thermal diffusivity of $D\approx 10^{-7}\ {\rm m^{2}/s}$.
@@
Note that $J_{d}=J_{h}^{2}$ when $v=Dd/h^{2}$, which, for the adhesive contribution where $h\approx 1\ {\rm nm}$, corresponds to an extremely high velocity in most cases. For example, if $d=1\ {\rm mm}$, then $v\approx 10^{8}\ {\rm m/s}$. Hence, for the adhesive contribution, the $J_{h}^{2}$ term can always be neglected.
In contrast, for the viscoelastic contribution, energy is dissipated in a volume element extending a distance $h\approx d$ into the rubber, so that $v=Dd/h^{2}\approx D/d\approx 0.1\ {\rm mm/s}$ for $d=1\ {\rm mm}$. Thus, for the viscoelastic case, the $J_{h}^{2}$ term cannot be neglected. However, in this case a more detailed analysis (as presented in Ref. [22]) is required, since viscoelastic energy dissipation decays rapidly with depth into the rubber.
Since $J_{d}=1$ corresponds to the sliding speed $v=D/d\approx 10^{-4}\ {\rm m/s}$ (using $d=1\ {\rm mm}$), and since for this velocity $\Delta T\approx qd/\kappa_{\rm th}=\mu\sigma vd/\kappa_{\rm th}\approx 0.1$-$1\ {\rm K}$ for $\sigma\approx 1$-$10\ {\rm MPa}$ (see below), we conclude that frictional heating is negligible at sliding speeds below $v\approx 10^{-4}\ {\rm m/s}$.
@@
$\Delta T\approx{qd\over\kappa_{\rm th}}(0.39J_{d})^{-1/2}\approx 1.6{\mu_{\rm cont}\sigma\over\kappa_{\rm th}}(vDd)^{1/2}$ (27)
@@
Figure 18: The influence of the flash temperature on (a) the sliding friction coefficient $\mu$ and (b) the relative contact area $A/A_{0}$. The green and red lines in (a) represent the viscoelastic contribution to the friction, $\mu_{\rm visc}$, and the total friction coefficient, $\mu=\mu_{\rm visc}+\mu_{\rm cont}$. The solid lines correspond to calculations with the flash temperature included, and the dashed lines to those without it. The results are for a tire tread compound on a concrete surface at a background temperature of $30^{\circ}{\rm C}$ and a nominal contact pressure of $0.3\ {\rm MPa}$.
Fig. 18 shows the influence of the flash temperature on the sliding friction coefficient and the area of real contact. The green and red lines in (a) represent the calculated viscoelastic contribution to the friction, $\mu_{\rm visc}$, and the total friction coefficient, $\mu=\mu_{\rm visc}+\mu_{\rm cont}$. The solid lines correspond to calculations with the flash temperature included, and the dashed lines to those without it. Fig. 18(b) shows the relative contact area $A/A_{0}$, with and without the flash temperature. Note that the flash temperature makes the rubber elastically softer, which results in an increase in the contact area with increasing sliding speed for $v>1\ {\rm m/s}$.
Figure 19: Flash temperature as a function of sliding speed, estimated using (27) with typical parameter values: $D=10^{-7}\ {\rm m^{2}/s}$, $\kappa_{\rm th}=0.5\ {\rm W/(K\cdot m)}$, $d=0.1\ {\rm mm}$, $\sigma=10^{7}\ {\rm Pa}$, and $\mu_{\rm cont}=0.5$.
@@
In our analysis of rubber friction on road surfaces, we have found that including surface roughness down to length scales where the cumulative rms slope $\xi$ is of order 1 yields results in good agreement with experimental data. This typically corresponds to a cut-off wavenumber of $q_{1}=2\pi/\lambda_{\rm min}\approx 10^{6}\ {\rm m}^{-1}$. For a further discussion of the large-wavenumber cut-off, see Ref. [58].
@@
Even if the exact value of $\lambda_{\rm min}$ is not known, the total friction coefficient can still be expressed as in (12), where the contact area contribution is written as $\mu_{\rm cont}=\sigma_{\rm eff}A/\sigma_{0}A_{0}$. Here, $A$ is the real contact area, calculated by including all surface roughness down to the assumed cut-off length. In this case the shear stress is an effective shear stress which could have contributions both from viscoelasticity and from the adhesive interaction in the area of real contact.
@@
Thus, in some studies, $A$ has been chosen as the nominal contact area $A_{0}$, in which case $\mu_{\rm visc}$ arises solely from macroscopic deformations of the contact region. This approach was employed in Ref. [59], where a rigid sphere was slid on human skin, a viscoelastic material. In that case, the nominal contact was Hertzian-like, and the macroscopic contact area had a distorted circular shape. Under such conditions, all asperity-induced viscoelastic energy dissipation is effectively captured within the definition of $\sigma_{\rm eff}$.
@@
Another important issue concerns the role of adhesion. Here, we are not referring to the adhesive interaction between rubber molecules and the substrate within the real contact area, which gives rise to the frictional shear stress $\sigma_{\rm f}$. Instead, we refer to two distinct adhesion-related mechanisms: (i) the increase in real contact area due to adhesion, and (ii) the energy dissipation from opening cracks at the edges of asperity contacts, occurring at length scales where adhesion is relevant.
@@
The contribution of adhesion to the enhancement of contact area was investigated by Plagge and Hentschke [54], while the contribution from opening cracks was studied by Le Gal et al. [60] and by Carbone et al. [38]. In Ref. [38] and [55] Carbone et al. also showed that adhesion increases the contact area. These mechanisms may significantly affect friction in very soft rubber compounds sliding on smooth surfaces. However, for rough substrate surfaces with roughness on many length scales, such as road surfaces, experimental evidence suggests that this form of adhesion is often negligible.